Grade 12th passAlgebra

Convert each part of the ratio to percentage:
1:2:5

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to find total parts by adding the given ratio = 1 + 2 + 5 = 8
1stpart = 1/8 = (1/8) × 100 %
= (100/8) %
= 12.5%
2ndpart = 2/8 = (2/8) × 100%
= (200/8)
= 25%
3rdpart = 5/8 = (5/8) × 100%
= (500/8)
= 62.5%

Thanks